OpenStudy (anonymous):

if u roll 3 dice what is the probability that at least one shows a "6"?

OpenStudy (anonymous):

the answer is 191/216. I just don't know how to get that

hero (hero):

sheesh

OpenStudy (anonymous):

thennn how you know the answer and you dont know how to get the answer that doesnt make any type of sense

OpenStudy (anonymous):

oh here? i can repeat what i said

OpenStudy (anonymous):

the probability that none of them are three is \[(\frac{5}{6})^3\]

OpenStudy (anonymous):

so the probability that 'at least one is a three' is \[1-(\frac{5}{6})^3\]

OpenStudy (anonymous):

that is how you always compute for "at least one". compute "none" then subtract from one

OpenStudy (anonymous):

i get \[\frac{91}{216}\] using a calculator if that is important

OpenStudy (anonymous):

@rriver if any steps are not clear let me know